# mHC: Manifold-Constrained Hyper-Connections

### v1

@@ 1 Introduction
where $\mathbf{x}_{l}$ and $\mathbf{x}_{l+1}$ denote the input and output of the $l$-th layer, respectively. Unlike the formulation in Eq. (1), the feature dimension of $\mathbf{x}_{l}$ and $\mathbf{x}_{l+1}$ is expanded from $C$ to $n\times C$, where $n$ is the expansion rate. The term $\mathcal{H}_{l}^{\mathrm{res}}\in\mathbb{R}^{n\times n}$ represents a learnable mapping that mixes features within the residual stream. Also as a learnable mapping, $\mathcal{H}_{l}^{\mathrm{pre}}\in\mathbb{R}^{1\times n}$ aggregates features from the nCnC-dim stream into a $C$-dim layer input, and conversely, $\mathcal{H}_{l}^{\mathrm{post}}\in\mathbb{R}^{1\times n}$ maps the layer output back onto the stream.
@@
To address these challenges, we propose Manifold-Constrained Hyper-Connections (mHC), as shown in Fig. 1(c), a general framework that projects the residual connection space of HC onto a specific manifold to restore the identity mapping property, while incorporating rigorous infrastructure optimization to ensure efficiency. Specifically, mHC utilizes the Sinkhorn-Knopp algorithm (sinkhorn1967concerning) to entropically project $\mathcal{H}_{l}^{\mathrm{res}}$ onto the Birkhoff polytope. This operation effectively constrains the residual connection matrices within the manifold that is constituted by doubly stochastic matrices. Since the row and column sums of these matrices equal to 11, the operation $\mathcal{H}_{l}^{\mathrm{res}}\mathbf{x}_{l}$ functions as a convex combination of the input features. This characteristic facilitates a well-conditioned signal propagation where the feature mean is conserved, and the signal norm is strictly regularized, effectively mitigating the risk of vanishing or exploding signals. Furthermore, due to the closure of matrix multiplication for doubly stochastic matrices, the composite mapping $\prod_{i=1}^{L-l}\mathcal{H}_{L-i}^{\mathrm{res}}$ retains this conservation property. Consequently, mHC effectively maintains the stability of identity mappings between arbitrary depths. To ensure efficiency, we employ kernel fusion and develop mixed precision kernels utilizing TileLang (wang2025tilelang). Furthermore, we mitigate the memory footprint through selective recomputing and carefully overlap communication within the DualPipe schedule (liu2024deepseek_v3).
@@ 4.1 Manifold-Constrained Hyper-Connections
It is worth noting that when $n=1$, the doubly stochastic condition degenerates to the scalar 11, thereby recovering the original identity mapping. The choice of double stochasticity confers several rigorous theoretical properties beneficial for large-scale model training:
@@ 4.3.1 Kernel Fusion
Using the coefficients derived from the aforementioned kernels, we introduce two additional kernels to apply these mappings: one for $\mathcal{F}_{\mathrm{pre}}\coloneq\mathcal{H}^{\mathrm{pre}}_{l}\mathbf{x}_{l}$ and another for $\mathcal{F}_{\mathrm{post,res}}\coloneq\mathcal{H}^{\mathrm{res}}_{l}\mathbf{x}_{l}+\mathcal{H}_{l}^{\mathrm{post}\,\top}\mathcal{F}(\cdot,\cdot)$. Through fusing the application of $\mathcal{H}^{\mathrm{post}}_{l}$ and $\mathcal{H}^{\mathrm{res}}_{l}$ with residual merging, we reduce the number of elements read from $(3n+1)C$ to $(n+1)C$ and the number of elements written from $3nC$ to nCnC for this kernel. We efficiently implement the majority of kernels (excluding Eq. (14) to (15)) using TileLang (wang2025tilelang). This framework streamlines the implementation of kernels with complex calculation process and allows us to fully utilize the memory bandwidth with minimal engineering effort.

### v2

@@ 1 Introduction
where $\mathbf{x}_{l}$ and $\mathbf{x}_{l+1}$ denote the input and output of the $l$-th layer, respectively. Unlike the formulation in Eq. (1), the feature dimension of $\mathbf{x}_{l}$ and $\mathbf{x}_{l+1}$ is expanded from $C$ to $n\times C$, where $n$ is the expansion rate. The term $\mathcal{H}_{l}^{\mathrm{res}}\in\mathbb{R}^{n\times n}$ represents a learnable mapping that mixes features within the residual stream. Also as a learnable mapping, $\mathcal{H}_{l}^{\mathrm{pre}}\in\mathbb{R}^{1\times n}$ aggregates features from the $nC$-dim stream into a $C$-dim layer input, and conversely, $\mathcal{H}_{l}^{\mathrm{post}}\in\mathbb{R}^{1\times n}$ maps the layer output back onto the stream.
@@
To address these challenges, we propose Manifold-Constrained Hyper-Connections (mHC), as shown in Fig. 1(c), a general framework that projects the residual connection space of HC onto a specific manifold to restore the identity mapping property, while incorporating rigorous infrastructure optimization to ensure efficiency. Specifically, mHC utilizes the Sinkhorn-Knopp algorithm (sinkhorn1967concerning) to entropically project $\mathcal{H}_{l}^{\mathrm{res}}$ onto the Birkhoff polytope. This operation effectively constrains the residual connection matrices within the manifold that is constituted by doubly stochastic matrices. Since the row and column sums of these matrices equal to $1$, the operation $\mathcal{H}_{l}^{\mathrm{res}}\mathbf{x}_{l}$ functions as a convex combination of the input features. This characteristic facilitates a well-conditioned signal propagation where the feature mean is conserved, and the signal norm is strictly regularized, effectively mitigating the risk of vanishing or exploding signals. Furthermore, due to the closure of matrix multiplication for doubly stochastic matrices, the composite mapping $\prod_{i=1}^{L-l}\mathcal{H}_{L-i}^{\mathrm{res}}$ retains this conservation property. Consequently, mHC effectively maintains the stability of identity mappings between arbitrary depths. To ensure efficiency, we employ kernel fusion and develop mixed precision kernels utilizing TileLang (wang2025tilelang). Furthermore, we mitigate the memory footprint through selective recomputing and carefully overlap communication within the DualPipe schedule (liu2024deepseek_v3).
@@ 4.1 Manifold-Constrained Hyper-Connections
It is worth noting that when $n=1$, the doubly stochastic condition degenerates to the scalar $1$, thereby recovering the original identity mapping. The choice of double stochasticity confers several rigorous theoretical properties beneficial for large-scale model training:
@@ 4.3.1 Kernel Fusion
Using the coefficients derived from the aforementioned kernels, we introduce two additional kernels to apply these mappings: one for $\mathcal{F}_{\mathrm{pre}}\coloneq\mathcal{H}^{\mathrm{pre}}_{l}\mathbf{x}_{l}$ and another for $\mathcal{F}_{\mathrm{post,res}}\coloneq\mathcal{H}^{\mathrm{res}}_{l}\mathbf{x}_{l}+\mathcal{H}_{l}^{\mathrm{post}\,\top}\mathcal{F}(\cdot,\cdot)$. Through fusing the application of $\mathcal{H}^{\mathrm{post}}_{l}$ and $\mathcal{H}^{\mathrm{res}}_{l}$ with residual merging, we reduce the number of elements read from $(3n+1)C$ to $(n+1)C$ and the number of elements written from $3nC$ to $nC$ for this kernel. We efficiently implement the majority of kernels (excluding Eq. (14) to (15)) using TileLang (wang2025tilelang). This framework streamlines the implementation of kernels with complex calculation process and allows us to fully utilize the memory bandwidth with minimal engineering effort.